# Mcity Data Engine: Iterative Model Improvement Through Open-Vocabulary Data Selection

### v1

@@ IV-A Selection of Open-Vocabulary Models
For a diverse selection, it is desirable to include a representative of each model architecture. For detections that include the classes of interest, a high recall is of interest, as it demonstrates a high true-positive ratio. While high precision is valuable for singular models, as it reduces false positives, it is of less importance in an ensemble, where model consensus reduces noise. While three models could not detect VRUs, the OWLv2 [25] family performs comparably well. Thus, the final model ensemble, as depicted in Fig. 2, consists of the five models OWLv2B, OWLv2L-E, OmDet-TurboT, Grounding-DINOT, and OWL-ViTL. Identifying VRU instances requires a $3/5$ majority consensus.
@@ IV-B Data Selection and Labeling
Figure 4: Data embeddings of frames (a) and VRUs (b) [32]: Orange dots represent data from $\mathcal{D}_{seed}$. Blue dots represent newly labeled data. Adding dark blue dots to $\mathcal{D}_{seed}$ forms $\mathcal{D}_{iter}$; adding all blue dots forms $\mathcal{D}_{iter}^{crowd}$, with light blue dots representing VRUs in crowds.
@@
Best mAP@0.5 $\mathcal{D}_{iter}$ 17.45% -0.22% 4.17% -2.45%
@@
Best F1 Score $\mathcal{D}_{iter}$ 12.37% 6.39% -0.84% 10.76%
@@
Best Recall $\mathcal{D}_{iter}$9.90% -12.35% -39.16% 26.05%
@@
Best Precision $\mathcal{D}_{iter}$ 14.95% -0.66% -8.37% 2.37%
@@
mAP@0.5 & F1 Score $\mathcal{D}_{iter}$ 14.50% 5.31% -6.08% 12.79%
@@
mAP@0.5 & Recall $\mathcal{D}_{iter}$ 15.03% -8.03% -22.58% 19.05%
@@ IV-C Iterative Model Improvement and Validation
Table III shows percentual improvements based on the prioritized metrics, including geometric means of multiple metrics. This model validation allows for an informed model selection for deployment. Based on the mAP@0.5 metric, the best results are achieved when crowd data is not included. Through a single iteration with the Mcity Data Engine, we can observe a significant improvement in mAP@50 of 17.45%, but this comes with slightly reduced recall. Training with the complete $\mathcal{D}_{iter}^{crowd}$ dataset shows an improvement in mAP@0.5 of only 6.91% but maintains recall compared to the model trained on $\mathcal{D}_{seed}$. If as few VRUs as possible should be missed, recall can be improved by up to 26.05%. Selecting model weights based on the geometric mean of both mAP@0.5 and F1 score can lead to significant improvements across most metrics. For the challenging use case of VRU detection in traffic data, we are interested in missing as few pedestrians as possible while not triggering too many false alarms. This suggests choosing model weights based on the $\mathcal{D}_{iter}$ training one at the highest geometric mean of mAP@0.5 and F1 score. Here, we see strong improvements in both mAP@0.5 and recall, with a slightly increased F1 score and a slight decrease in precision. Selecting model weights at the best recall value or at the best geometric mean of mAP@0.5 and recall leads to a drastic increase in false positives. The selected model weights can now be deployed in the cloud or exported and deployed on the edge.
@@ V Conclusion
In this work, we have first provided an overview of current data engines and have highlighted a lack of coverage and availability among academic data engines, as shown in Table I. Addressing this gap, the open-source Mcity Data Engine provides all modules necessary for the complete data-based development cycle, similar to full-fledged industrial proprietary solutions, as shown in Fig. 1. This enables researchers and the open-source community to leverage existing data sources and iteratively improve models.
By selecting data of interest with a robust ensemble-based open-vocabulary approach, rare and novel classes can be found in an efficient manner within large amounts of unlabeled data through the use of natural language, making the Mcity Data Engine accessible to non-experts. This approach is shown in Fig. 2.
For future work, we are interested in leveraging a wider selection of VLMs for data selection based on natural language. VLMs with improved prediction quality might reduce the need for a computationally intensive ensemble. In addition, we would like to evaluate our approach with more iterations of data labeling and model training to further improve the observed precision-recall tradeoff.
@@ Acknowledgment
This work was made possible through a scholarship provided by Mcity, funding Daniel Bogdoll as a visiting research scholar. The stay was supported by the KIT Research Travel Grant and the KIT Graduate School UpGrade Mobility, providing additional financial aid.

### v2

@@ IV-A Selection of Open-Vocabulary Models
For a diverse selection, it is desirable to include a representative of each model architecture. For detections that include the classes of interest, a high recall is of interest, as it demonstrates a high true-positive ratio. While high precision is valuable for singular models, as it reduces false positives, it is of less importance in an ensemble, where model consensus reduces noise. While three models could not detect VRUs, the OWLv2 [25] family performs comparably well. Thus, the final model ensemble, as depicted in Fig. 2, consists of the five models OWLv2B, OWLv2L-E, OmDet-TurboT, Grounding-DINOT, and OWL-ViTL. Identifying VRU instances requires a 3/53/53 / 5 majority consensus.
@@ IV-B Data Selection and Labeling
Figure 4: Data embeddings [32] of whole frames (a) and individual VRUs (b): Orange dots represent data from $\mathcal{D}_{seed}$. Blue dots represent newly labeled data. Adding dark blue dots to $\mathcal{D}_{seed}$ forms $\mathcal{D}_{iter}$; adding all blue dots forms $\mathcal{D}_{iter}^{crowd}$, with light blue dots representing VRUs in crowds.
@@
Best
mAP@0.5
$\mathcal{D}_{iter}$ 17.45% -0.22% 4.17% -2.45%
@@
Best
F1 Score
$\mathcal{D}_{iter}$ 12.37% 6.39% -0.84% 10.76%
@@
Best
Recall
$\mathcal{D}_{iter}$9.90% -12.35% -39.16% 26.05%
@@
Best
Precision
$\mathcal{D}_{iter}$ 14.95% -0.66% -8.37% 2.37%
@@
mAP@0.5 &
F1 Score
$\mathcal{D}_{iter}$ 14.50% 5.31% -6.08% 12.79%
@@
mAP@0.5 &
Recall
$\mathcal{D}_{iter}$ 15.03% -8.03% -22.58% 19.05%
@@ IV-C Iterative Model Improvement and Validation
Table III shows percentual improvements based on the prioritized metrics, including geometric means of multiple metrics. This model validation allows for an informed model selection for deployment. Based on the mAP@0.5 metric, the best results are achieved when crowd data is not included. Through a single iteration with the Mcity Data Engine, we can observe a significant improvement in mAP@0.5 of 17.45%, but this comes with slightly reduced recall. Training with the complete $\mathcal{D}_{iter}^{crowd}$ dataset shows an improvement in mAP@0.5 of only 6.91% but maintains recall compared to the model trained on $\mathcal{D}_{seed}$. If as few VRUs as possible should be missed, recall can be improved by up to 26.05%. Selecting model weights based on the geometric mean of both mAP@0.5 and F1 score can lead to significant improvements across most metrics. For the challenging use case of VRU detection in traffic data, we are interested in missing as few pedestrians as possible while not triggering too many false alarms. This suggests choosing model weights based on the $\mathcal{D}_{iter}$ training one at the highest geometric mean of mAP@0.5 and F1 score. Here, we see strong improvements in both mAP@0.5 and recall, with a slightly increased F1 score and a slight decrease in precision. Selecting model weights at the best recall value or at the best geometric mean of mAP@0.5 and recall leads to a drastic increase in false positives. The selected model weights can now be deployed in the cloud or exported and deployed on the edge.
@@ V Conclusion
In this work, we have first provided an overview of current data engines and have highlighted a lack of coverage and availability among academic data engines, as shown in Table I. Addressing this gap, the open-source Mcity Data Engine provides all modules necessary for the complete data-based development cycle, similar to full-fledged industrial proprietary solutions, as shown in Fig. 1. This enables researchers and the open-source community to leverage existing data sources and iteratively improve models without the need to develop their own solutions while remaining flexible enough to extend and adapt the MDE based on their needs.
By selecting data of interest with a robust ensemble-based open-vocabulary approach, rare and novel classes can be found in an efficient manner within large amounts of unlabeled data through the use of natural language, making the MDE accessible to non-experts as shown in Fig. 2.
For future work, we are interested in leveraging a wider selection of VLMs for data selection based on natural language. VLMs with improved prediction quality might reduce the need for a computationally intensive ensemble. In addition, we would like to evaluate our approach with more iterations of data labeling and model training to further improve the observed precision-recall tradeoff. Finally, integrating data types beyond 2D images into the MDE is of interest to us.
@@ Acknowledgment
This work was made possible through a scholarship provided by Mcity, funding Daniel Bogdoll as a visiting research scholar. The stay was financially supported by the KIT Research Travel Grant and the KIT Graduate School UpGrade Mobility.